OpenStudy (anonymous):

-5y -9 = -(y-1) Answer Choices: -1/2 -2 1/2 -2 -2/5 Thank You!! :]

OpenStudy (anonymous):

-5y-9=-(y-1) -5y-9=-y+1 +5y +5y -9=4y+1 -1 -1 -10=4y y= -10/4 -5/2 so the answer is b

OpenStudy (anonymous):

Thank you! [:

OpenStudy (anonymous):

wait.. but b is 2 and one half :o

OpenStudy (anonymous):

So how is it b?

OpenStudy (anonymous):

Can you tell me @Dzie ?

OpenStudy (anonymous):

you can reduce the -5/2 to a mixed fraction that is -5 divided by 2 to get -2half which is the second answer

OpenStudy (anonymous):

So that is the final answer? sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you for explaining! :]

OpenStudy (anonymous):

it is my pleasure

OpenStudy (anonymous):

i hope you are okay with the explanation